Solve for k in the equation -8 + 44 = -20 + 80 and explain each step.


1
Expert's answer
2021-06-01T14:06:11-0400

"-8k+44=-20k+80"

Simplify by combining like terms

"-8k+20k=80-44"

"12k=36"

Divide both sides by 12 to get value of k


"\frac{12k}{12}=\frac{36}{12}"


"K=3"